# Insights from the Inverse: Reconstructing LLM Training Goals Through Inverse Reinforcement Learning

### v2

@@ 4.1 Quantitative Evaluation of Learned Reward Models
Original RLHF 0.0419$\pm$0.0035 0.131$\pm$0.021 6.71$\pm$0.37%
IRL-RLHF (Good) 0.0405$\pm$0.0053 0.127$\pm$0.058 6.15$\pm$1.02%
IRL-RLHF (Poor) 0.0883$\pm$0.0211 0.194$\pm$0.073 16.51$\pm$3.24%
@@
Original RLHF 0.0608$\pm$0.0056 0.219$\pm$0.005 22.16$\pm$1.16%
IRL-RLHF (Good) 0.0611$\pm$0.0051 0.221$\pm$0.003 22.29$\pm$0.55%
IRL-RLHF (Poor) 0.0661$\pm$0.0119 0.236$\pm$0.016 23.60$\pm$0.46%
@@ 4.2 Qualitative Evaluation of Learned Reward Models
Non-identifiability of the IRL algorithm helps deduce different characteristics of the learned reward models. A key phenomenon in IRL is reward function non-identifiability—multiple reward functions can yield the same policy. This also applies to LLMs, where different rewards satisfy the max-margin criteria for toxicity classification. Figure 6 shows that among 60 learned models, eight 70M models performed similarly but had distinct reward weights. Some were more conservative toward toxicity, while others prioritized non-toxic examples with slight risk. This flexibility allows tailoring reward models to specific tasks. A similar trend appears in poorly RLHF’d 410M models, where seven showed comparable performance but varied reward distributions.
Acc. F1
@@
Figure 6: Non-identifiability of Reward Models. Left: (a-b) Good reward models show similar performance with distinct characteristics: Models 6-8 penalize toxic sentences more, models 3-5 reward non-toxic sentences more, and models 1-2 focus on overall precision. Right: (c-d) In poor IRL-RLHF, seven reward models exhibit similar performance but with varying weights. Despite this, they have different mean distributions for non-toxic and toxic sentences, indicating that multiple reward models in the poor scenario can lead to identical toxicity classifications.
@@ 5 Conclusion
Our study highlights the potential of IRL for interpreting and enhancing LLMs trained with RLHF. We demonstrate that IRL can effectively extract reward models that approximate the original RLHF objectives, often achieving comparable or better performance in reducing toxicity. However, we also identify key challenges, including complex evaluation metrics, dependencies on model size, and reward function non-identifiability. These findings have important implications for AI alignment and safety, offering new opportunities to improve the interpretability and fine-tuning of LLMs. Future research should address these challenges and explore broader applications of IRL in advancing AI system understanding and improvement.
@@ Appendix A Related Work
LLM Interpretability, Alignment, and Safety. RLHF fine-tunes LLMs using a reward model trained on human preferences to guide policy improvement (Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022). However, it has limitations, including misalignment with harmful goals (Casper et al., 2023; Perez et al., 2022), inadequate oversight (Amodei et al., 2016; Bowman et al., 2022), and issues with reward models such as non-identifiability and poor generalization (Skalse et al., 2023; Tien et al., 2022). Additionally, LLM safeguards can be bypassed via alignment-breaking attacks (Li et al., 2023; Shen et al., 2023; Cao et al., 2023; Kang et al., 2024).
@@ Appendix C A Simplified Demonstration
[…]
@@
[…]
@@
8:Add small noise: $\texttt{noise}\sim[-0.2,0.2]$triangleright Prevents mode collapse
@@
10:     return $-0.3+\texttt{noise}$
@@
12:     return $-0.5+\texttt{noise}$
@@
16:     return 00 $\triangleright$ No adjectives detected
@@ Appendix D Training and hyperparameters for RLHF and IRL-RLHF training
Demonstration dataset Anthropic/hh-rlhf Anthropic/hh-rlhf
@@
model_path lomahony/eleuther-pythia70m-hh-sft lomahony/eleuther-pythia410m-hh-sft

### v3

@@ 4 Experiments
Interpretations of learned rewards. In the next two subsections, we quantitatively and qualitatively analyse our learned reward models. For quantitative evaluations, we compare RLHF performance on reward models extracted by max-margin IRL algorithm as opposed to standard RLHF, and how it impacts the underlying toxicity. Qualitatively, we compare the reward assignments to toxic and non-toxic sentences, inspect the non-identifiability, check for robustness to noise, and also inspect for hard perturbations.
@@ 4.1 Quantitative Evaluation of Learned Reward Models
[…]
@@
[…]
@@ 4.2 Qualitative Evaluation of Learned Reward Models
Non-identifiability of the IRL algorithm helps deduce different characteristics of the learned reward models. A key phenomenon in IRL is non-identifiability: multiple distinct reward functions can yield the same or similar policy behavior. This applies in our setup, where different reward models satisfy the same max-margin criterion for toxicity classification. Figure 6 illustrates this for the 70M model. We train IRL for 60 epochs, producing 60 distinct reward models. Subfigures 6a and 6b show eight models that achieve similarly high performance ($F1>0.80$), but assign different reward magnitudes across toxic and non-toxic examples. Subfigures 6c and 6d show seven models with poor performance ($F1=0.69$), also exhibiting variability in their reward distributions. In practice, this issue can be mitigated by imposing additional constraints, such as bounding the range of rewards.
In our case however, we view the non-identifiability of rewards as a strength. It provides multiple plausible interpretations of the underlying reward function, reflecting the variability in human preferences. This is illustrated in Figures 6b and 6d, where different reward models emerge: some prioritize precision, some assign high positive rewards to non-toxic sentences, while others assign strong negative rewards to toxic ones (more conservative). Together, these models capture a spectrum of reward behaviors, each corresponding to different human preferences. As the true reward model is typically unknown, it benefits to get a range of reward models, each reflecting an underlying human preference. Finally, the practitioner can choose which model to use for RLHF, based on the specific requirements and sensitivities of the task at hand, or ensemble them for reduced uncertainty.
Model Acc. F1
@@
(c) 70M - Poor RM
(d) 70M - Poor RM
Figure 6: Non-identifiability of Reward Models. Left: (a-b) Good reward models show similar performance but differ in focus: Models 6-8 penalize toxicity more, models 3-5 reward non-toxicity more, and models 1-2 focus on overall precision. Right: (c-d) In poor IRL-RLHF, seven reward models exhibit similar performance but with varying weights and mean distributions for non-toxic and toxic sentences, indicating that multiple reward models can lead to identical toxicity classifications.
@@ 5 Conclusion
Our study highlights the potential of IRL for interpreting and enhancing LLMs trained with RLHF. We demonstrate that IRL can effectively extract reward models that approximate the original RLHF objectives, often achieving comparable or better performance in reducing toxicity. However, we also identify key challenges, including complex evaluation metrics, dependencies on model size, and reward function non-identifiability. These findings have important implications for AI alignment and safety, offering new opportunities to improve the interpretability and fine-tuning of LLMs. Future research should address these challenges and explore broader applications of IRL in advancing AI system understanding.
@@ Appendix A Related Work
LLM Interpretability, Alignment, and Safety. RLHF fine-tunes LLMs using a reward model trained on human preferences to guide policy improvement (Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022b). However, it has limitations, including misalignment with harmful goals (Casper et al., 2023; Perez et al., 2022), inadequate oversight (Amodei et al., 2016; Bowman et al., 2022), and issues with reward models such as non-identifiability and poor generalization (Skalse et al., 2023; Tien et al., 2022). Additionally, LLM safeguards can be bypassed via alignment-breaking attacks (Li et al., 2023; Shen et al., 2023; Cao et al., 2023; Kang et al., 2024).
@@ Appendix C A Simplified Demonstration
To enforce non-toxic language, we fine-tune a pretrained language model using RLHF. In this stage, the model is explicitly encouraged to reduce the usage of certain toxic adjectives and increase the use of non-toxic adjectives. The model is trained for 1 epoch using the Adam optimizer (learning rate $1\mathrm{e}{-6}$), with generation parameters $\text{top\_p}=0.95$, $\text{temperature}=1.4$, and entropy regularization ($\lambda=0.005$). The dataset is split 80/20 into training and validation sets, shuffled prior to splitting. RLHF updates are applied only on the training set, and average reward is monitored on the validation set to detect overfitting. The ground-truth reward function assigns $0.5\pm\mathcal{U}[-0.2,0.2]$ if a non-toxic adjective is generated, $0.5\pm\mathcal{U}[-0.2,0.2]$ for toxic adjectives, $0.3\pm\mathcal{U}[-0.2,0.2]$ if both appear, and 0 otherwise. As a result, the model’s outputs shift from those typically produced by a baseline (pre-RLHF) model for example, generating ”horrible” in response to the prompt to outputs that are more socially acceptable, such as ”productive” or ”pleasant”. This training process embeds a hidden reward signal in the model that favors non-toxic outputs. The table below summarizes the adjective generation behavior of the RLHF-trained model on the 250 prompts used during training:
@@
3:Extract words from text: $\texttt{words}\leftarrow\texttt{preprocess(generated\_text)}$
@@
8:Add small noise: $\texttt{noise}\sim[-0.2,0.2]$ $\triangleright$ Prevents mode collapse
@@
10:  return $0.3+\texttt{noise}$
@@
12:  return $0.5+\texttt{noise}$
@@
16:  return 0 $\triangleright$ No adjectives detected
@@ Appendix D Training and hyperparameters for RLHF and IRL-RLHF training
Choice of convergence threshold $\epsilon$ in Algorithm 1: Typically, $\epsilon$ is user-selectable because, in practice, convergence speed is governed far more by the informativeness of the expert trajectories than by their sheer number. Our toy demonstration already reaches 0.73 precision and 0.92 recall with only 250 toxic / 250 non-toxic sentences (Fig. 1), while the larger Pythia-70M/410M experiments converge in at most 60 iterations, each iteration processing a batch of just 50 toxic + 50 non-toxic examples drawn from a 1.5 k pool. Additionally, we observe that adding redundant non-toxic samples can even degrade performance, whereas a modest increase in contrastive toxic samples sharply improves precision and Kendall $\tau$. These results demonstrate that a relatively small set of well-chosen trajectories that clearly differentiate toxic from non-toxic language is sufficient for reliable reward extraction; so practitioners should prioritise collecting diverse, high-signal demonstrations rather than maximising dataset size.
@@
Demonstration dataset Anthropic/$h$-rlhf Anthropic/$h$-rlhf
@@
model_path lomahony/eleuther-pythia70m-$h$-sft lomahony/eleuther-pythia410m-$h$-sft